OpenStudy (anonymous):

In your lab, a substance's temperature has been observed to follow the function T(x) = (x - 4)3 + 6. The turning point of the graph is where the substance changes from a liquid to a gas. Explain to your fellow scientists how to find the turning point of this function, using complete sentences.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

Do you know about the turning point formula?

OpenStudy (anonymous):

no but since i read your attachment i do now @phi

OpenStudy (anonymous):

so A= 4 B=3 C=6 right

OpenStudy (phi):

I think the a,b, and c are from a x^3 + b x^2 +c x + d unfortunately, you have (x - 4)^3 + 6 which is (x-4) (x-4) (x-4) + 6 we have to multiply that out to get it in the form so we can find a, b and c

OpenStudy (anonymous):

so to multiple you have distribute

OpenStudy (phi):

yes. but in steps first do (x-4)(x-4)

OpenStudy (anonymous):

\[x ^{2}-8x +16\]

OpenStudy (phi):

now do (x^2 -8x+16)(x-4)

OpenStudy (anonymous):

\[x ^{3}-12x ^{2}+16x-64\]

OpenStudy (phi):

you missed a +32x

OpenStudy (anonymous):

oh hold im redoing it

OpenStudy (anonymous):

\[x ^{3}-12x ^{2}+48x-64\]

OpenStudy (phi):

that means you function is \[ (x - 4)^ 3 + 6 = x ^{3}-12x ^{2}+48x-64 + 6 \\ =x ^{3}-12x ^{2}+48x - 58 \]

OpenStudy (anonymous):

ok so now does that means that A=12 B=48 C=-58

OpenStudy (phi):

almost, you start with x^3 (read that as 1*x^3... so A=1, and so on)

OpenStudy (anonymous):

oh i didnt see the x^3

OpenStudy (anonymous):

so A=1 B=12 C=48

OpenStudy (phi):

even closer, but include the signs

OpenStudy (anonymous):

yea my bad B=-12

OpenStudy (phi):

now we use the formula

OpenStudy (anonymous):

ok i had written the formula down on a piece of paper

OpenStudy (anonymous):

\[-12\sqrt{0} \over 3\]

OpenStudy (phi):

the formula says -B in front. B is -12, so you have - -12 or +12

OpenStudy (phi):

and you are adding the sqr(0) so it is \[ \frac{12±\sqrt{0}}{3} \] sqrt(0) is 0 so that is \[ \frac{12±0}{3} \\\frac{12}{3} \]

OpenStudy (phi):

the answer is x=12/3 = 4 that is the turning point of this curve Here is a graph

OpenStudy (anonymous):

ok thank you so much for your help